# Shaping Shared Languages: Human and Large Language Models’ Inductive Biases in Emergent Communication

### v1

@@ Abstract
Languages are shaped by the inductive biases of their users. Using a classical referential game, we investigate how artificial languages evolve when optimised for inductive biases in humans and large language models (LLMs) via Human-Human, LLM-LLM and Human-LLM experiments. We show that referentially grounded vocabularies emerge that enable reliable communication in all conditions, even when humans and LLMs collaborate. Comparisons between conditions reveal that languages optimised for LLMs subtly differ from those optimised for humans. Interestingly, interactions between humans and LLMs alleviate these differences and result in vocabularies which are more human-like than LLM-like. These findings advance our understanding of how inductive biases in LLMs play a role in the dynamic nature of human language and contribute to maintaining alignment in human and machine communication. In particular, our work underscores the need to think of new methods that include human interaction in the training processes of LLMs, and shows that using communicative success as a reward signal can be a fruitful, novel direction.
@@ 1 Introduction
Languages adapt to how they are learned and used. The primary reason is the continuous influence of individuals’ (learning) biases and pressures that slowly shape languages to become more structured, easier to learn and communicatively efficient Smith (2022). Although a wealth of experiments in the field of language evolution has contributed to this knowledge (Kirby et al., 2014; Raviv et al., 2019, i.a.), only relatively recently we started investigating whether these principles can be applied to large language models as well Galke et al. (2024); Kouwenhoven et al. (2025). For instance, more systematic and structured languages are typically easier to learn by humans who are asked to learn novel artificial languages Raviv et al. (2021). Recent work by Galke et al. (2024) revealed that the same is true for recurrent neural networks and LLMs. Moreover, transmission of initially unstructured language systems over generations of human learners (i.e., iterated learning) increases structure and learnability in these languages Kirby et al. (2015). To investigate whether this process leads to a similar outcome with LLMs, Kouwenhoven et al. (2025) created a setting in which LLMs learned an initially holistic unstructured artificial language and then used this repeatedly to communicate in a referential game. They showed that the linguistic structure of these languages increased, which enabled more successful communication between LLM agents, again mirroring observations from human experiments Kirby et al. (2015).
@@
With AI systems being increasingly incorporated into our daily lives Brinkmann et al. (2023), it is argued that repeated interactions with machines become increasingly important to maintain alignment Mikolov et al. (2018); Beuls and Van Eecke (2024) and referential grounding Kouwenhoven et al. (2022). In the case of humans, these repeated interactions cause languages to evolve such that they accommodate for specific abilities and preferences of minority individuals at the group level Josserand et al. (2024). Since the seemingly similar ways that languages adapt and optimise as a result of learning and use in both, Human-Human and LLM-LLM interactions, the question that arises is whether these processes can also be used to experimentally evolve a language that is optimised for humans and LLMs. In other words; can humans and LLMs collaboratively shape a language that is easy to learn for both, and allows for successful communication? If so, what do these languages look like?
@@
This is investigated here by extending earlier work by Kouwenhoven et al. (2025). Firstly, we provide experimental data of humans playing the same referential game that Kouwenhoven et al. (2025) used with LLMs, allowing comparisons between languages evolved through LLM-LLM interactions with those resulting from Human-Human interactions. Secondly, we run experiments where humans collaborate with an LLM (Figure 1)111This study was approved by the ethics department of Leiden University (2024-03-11-R.E. de Kleijn-V1-5354). While unclear how human and LLM abilities exactly differ, this allows us to test whether an artificial holistic language can be optimised for the inductive biases of two different types of language learners. If shared vocabularies of signals and meanings emerge that allow for successful communication, one could argue that there has been some form of referential grounding, a prerequisite for successful communication Clark and Brennan (1991). Finally, Human-LLM collaboration allows investigating if and how the evolved language differ from languages that evolve within Human-Human and LLM-LLM interactions.
@@
Our results show that structured and referentially grounded languages can emerge when humans and LLMs interact repeatedly in our experiment. The emergent languages from these interactions tend to be more human-like than LLM-like, suggesting that the LLMs are flexible towards the strong human preferences that shape the languages. Finally, languages optimised for LLMs result in less variation and are more degenerate than those optimised for humans.
@@ 2 Background
In this section we discuss relevant work in the field of language evolution, the role of inductive biases in language evolution, and why this is relevant for LLM research.
@@ 2.1 Language evolution
Language allows us to communicate successfully because of the vocabulary we share but also due to its open-ended nature, enabling the possibility to express novel meanings via compositional semantics. This defining feature of human language means that the meaning of any phrase is derived from the meanings of its individual components and the rules by which they are combined Hockett (1960). The evolution of compositionality has been investigated abundantly in the field of language evolution through human experiments (Kirby et al., 2008; Raviv et al., 2019, e.g.) and computational simulations (de Boer, 2006; Steels et al., 2012; Lazaridou and Baroni, 2020; Kouwenhoven et al., 2024, e.g.). These experiments typically involve learning artificial languages or playing a signalling game. Here, learning artificial languages imposes a constraint for which it is believed to lead to more structured languages Kirby et al. (2015). Communication in signalling games imposes a pressure for expressivity, requiring participants to develop a vocabulary of signal-meaning mappings that allows them to communicate about novel stimuli. In this case, some form of referential grounding must be established through the process of repeated interactions. Participants–human or machine–generally establish novel signal-meaning mappings quickly that enable successful communication.
@@ 2.2 Inductive biases
An important aspect in this paper is the notion of biases. Here, we do not so much focus on behavioural biases observed in humans (e.g., the confirmation bias), but are rather interested in implicit inductive biases that may result in biased language learning. This is relevant since seemingly arbitrary aspects of linguistic structure may actually result from general learning and processing biases deriving from the structure of thought processes, perceptuo-motor factors, cognitive limitations, and pragmatics Christiansen and Chater (2008). For humans, inductive biases like a preference for compressibility, learnability, simplicity, or efficiency Kirby et al. (2015); Tamariz and Kirby (2015); Gibson et al. (2019) naturally shape how languages evolve. Some even argue that humans’ cognitive limitations may be beneficial for language acquisition DeCaro et al. (2008); Poletiek et al. (2018). In the case of simulations of (reinforcement learning) agents, inductive biases typically do not match those present in humans. As such, they are often induced artificially by incorporating biases to guide learning dynamics as a means to recover human-like properties (Galke and Raviv, 2025, for a review see). In the case of LLMs, being fundamentally different from humans (although this is an ongoing debate Kozachkov et al. (2023)), we focus on increasingly apparent inductive biases of the transformer architecture Futrell and Mahowald (2025) that may influence how languages evolve in the context of our experiment.
@@
Although the underlying mechanisms of these biases differ between humans and machines, we find substantial overlap in terms of their behavioural effects. As such we hypothesize that the aforementioned effects of continuous learning and use of language will also come into play when humans and machines collaborate and result in a language optimised for the preferences of both entities.
@@ 3 Methodology
This experiment revolves around the classical referential Lewis game as implemented in Kouwenhoven et al. (2025), who based their setup on previous work in emergent communication(Raviv et al., 2021; Kirby et al., 2015, e.g.). We extend the setup used by Kouwenhoven et al. (2025) to incorporate humans. In total, 45 participants participated in the experiment, 30 of which formed 15 Human-Human pairs, and the remaining 15 interacted with an LLM in a Human-LLM setup. This allows us to directly compare languages adapted for human preferences to those adapted for the LLM-LLM simulations Kouwenhoven et al. (2025). But perhaps most interestingly, the Human-LLM condition provides an opportunity to investigate whether languages can be optimised for entities with different mechanisms and cognitive capacities (e.g., memory), and if so, we can see what these look like.
During the experiment, participants learn an artificial language and use it during communication. This language comprises a meaning space consisting of three attributes (shape, colour, and amount) that each can have three values, totalling to 27 unique stimuli. The corresponding labels are initialised following the design of Kirby et al. (2008), creating a holistic artificial language without structure (e.g., ”watopo”, ”sanasowi”, ”pikuku”) that contained a limited set of characters to prevent participants from writing English words. Participants first learn 15 random signal-meaning pairs individually through the exposure, guessing, and labelling blocks. Hereafter, participants use the newly acquired language to communicate in a referential game. In this game, the speaker observes a target stimulus and labels it. Using this label as a signal, the listener is then tasked with identifying the correct target among three distractors. Cooperation is successful when the listeners’ guess matches the target. After the communication block, there is a testing block in which participants individually label 27 meanings, including 12 unseen meanings. The duration of the entire experiment is roughly 70 minutes. An experiment overview is provided in Figure 1.
@@ 3.1 LLMs as participants
Human participants learn the language by going through the exposure and guessing blocks twice. They iteratively go over the 15 training stimuli and extract some apparent, but not present, patterns or consistencies. They are then tasked to label the stimuli, before moving on to the communication block in which they interact with another human or a LLM. In either case, they were told that they interacted with a human. The LLM agents, however, are not updated and receive instructions to learn the languages by virtue of in-context learning. Specifically, the stimuli are presented in a structured JSON-like format (Figure 2) that proved effective in Galke et al. (2024) and Kouwenhoven et al. (2025). As such, we assume that these signal-meaning mappings in the context of a prompt provide enough (distributional) information for a LLM to learn a mapping between the attributes of the stimuli and signal syllables (1). Although the prompt structure ‘invites’ the LLM to infer a signal from the stimulus attributes, we are agnostic about how exactly and what kind of mapping is deduced but are interested in the resulting behaviours. In the experiments, we use the instruction-tuned variant of Llama 3 70B Llama Team (2024) with greedy sampling. We opt for an instruction-tuned model since this allows us to specify the need for communicative success. This potentially affects how the model’s inductive biases are expressed, but we leave this to future work.
@@
One of essentially two tasks is performed throughout the experiment: labelling or guessing. The labelling block and speaking in the communication block involve labelling, and the guessing block and discrimination during communication involve guessing. Generating signals is achieved through prompt completion. Guessing is done by prefilling the prompt with distractor stimuli or labels and selecting the item with the highest probability. This alleviates LLMs’ inconsistent behaviour in answering multiple-choice questions Khatun and Brown (2024), and follows recommendations from computational linguistics Hendrycks et al. (2021); Wang et al. (2024). During communication, we add a communicativeSuccess attribute set to if the previous interaction for this stimulus was successful and zero otherwise. This attribute functions as a memory between interactions and acts as a pressure for expressivity. In human language evolution, such pressure plays an important role since it prevents languages from becoming degenerate Kirby et al. (2015). Importantly, the agents observe the training vocabulary in their context with the current stimulus in the guessing and labelling block, rendering them as simple look-up tasks. We do, however, not include the current stimulus during communication and testing, requiring the agents to extract an appropriate mapping and generalise to new stimuli. Akin to common practise in older simulations (Steels et al., 2012, e.g.)), the agent vocabularies are updated when labels are generated after the labelling block and during the communication block. This allows the vocabularies of signal-meaning mappings to evolve during the simulation. As such, prompts are slightly different after each interaction. Moreover, given the primacy and recency bias in LLMs Liu et al. (2024); Mina et al. (2025), we shuffle the vocabulary before creating prompts to account for unwanted ordering effects.
@@ 3.2 Metrics
Besides comparing the percentage of communicative success (PercCom), the primary goal of this work is to understand what a language looks like when optimised for different entities. Specifically, we investigate whether the languages display some degree of structure in the form of compositionality. In this experiment, this means that attribute values are denoted with label parts that are re-used to describe other similar stimuli. Capturing this is not at all trivial, especially provided the freedom given to participants when they label stimuli. A common metric that gauges whether similar meanings map to similar signals is Topographic Similarity (Brighton and Kirby, 2006, TopSim). While providing a good indication of compositional language use, it does not take variability in language, such as word-order freedom, into account. It could therefore show an incomplete picture (i.e., a low TopSim) since languages can still be compositional despite having multiple word orders, the existence of synonyms, or homonyms Conklin and Smith (2023). Hence, we report multiple metrics in addition to TopSim that together indicate the degree of compositionality. Specifically, on synonymy (one-to-many mappings), homonymy (many-to-one), and word order freedom (Freedom), for which Conklin and Smith (2023) proposed entropy-based metrics. A language where each attribute value is encoded by a single character in a position has low entropy, and thus a low synonymy. Languages with a uniform distribution over all characters to refer to an attribute value have high synonymy. Homonymy is similar, it looks at how many attribute-values a character in a position can refer to, i.e., when characters can map to multiple attribute-values. Finally, we compute word-order freedom (freedom) to account for variability in the order by which labels are composed. It assesses whether each value of a specific attribute is encoded in a specific position of the label, i.e., there is little freedom, or whether attribute values can be encoded in any position of the label, i.e., displaying a high degree of word order freedom ().
@@
Systematic generalisation to novel stimuli is assessed through the generalisation score GenScore from Raviv et al. (2021). It gauges whether the labels produced for unknown (i.e., testing) stimuli are labelled in consistent ways to labels produced for similar known (i.e., training) stimuli. In addition to character-based metrics, we assess whether participants re-use parts of labels in different labels by computing the Ngram diversity Li et al. (2016) over all the produced labels in a block. Ngram diversity is the average ratio of unique vs. total Ngrams for in all labels. Low Ngram diversity implies that labels are composed of re-used parts and high diversity means that labels do not share many Ngrams, thus are very different. The percentage of unique labels captures the degree of degeneracy (RatioUniLabels). Finally, we measure whether a pressure for communicative success, known to drive efficiency in human experiments Smith and Culbertson (2020), results in shorter labels using WordLength.

### v2

@@ Abstract
Languages are shaped by the inductive biases of their users. Using a classical referential game, we investigate how artificial languages evolve when optimised for inductive biases in humans and large language models (LLMs) via Human-Human, LLM-LLM and Human-LLM experiments. We show that referentially grounded vocabularies emerge that enable reliable communication in all conditions, even when humans and LLMs collaborate. Comparisons between conditions reveal that languages optimised for LLMs subtly differ from those optimised for humans. Interestingly, interactions between humans and LLMs alleviate these differences and result in vocabularies more human-like than LLM-like. These findings advance our understanding of the role inductive biases in LLMs play in the dynamic nature of human language and contribute to maintaining alignment in human and machine communication. In particular, our work underscores the need to think of new LLM training methods that include human interaction and shows that using communicative success as a reward signal can be a fruitful, novel direction.
@@ 1 Introduction
Languages adapt to how they are learned and used. The primary reason is the continuous influence of individuals’ (learning) biases and pressures that slowly shape languages to become more structured, easier to learn and communicatively efficient Smith (2022). Although a wealth of experiments in the field of language evolution have contributed to this (Kirby et al., 2014; Raviv et al., 2019, i.a.), only relatively recently have we started investigating whether these principles can be applied to large language models as well Galke et al. (2024); Kouwenhoven et al. (2025). For instance, more systematic and structured languages are typically easier to learn by humans who are asked to learn novel artificial languages Raviv et al. (2021). Recent work by Galke et al. (2024) showed that the same is true for recurrent neural networks and LLMs. Moreover, transmission of initially unstructured language systems over generations of human learners (i.e., iterated learning) increases structure and learnability in these languages Kirby et al. (2015). To investigate if LLMs show similar outcomes, Kouwenhoven et al. (2025) created a setting in which LLMs learned an initially holistic unstructured artificial language and then used this repeatedly to communicate in a referential game. They showed that the linguistic structure of these languages increased and more successful communication between LLM agents—again mirroring observations from human experiments Kirby et al. (2015).
@@
With AI systems being increasingly incorporated into our daily lives Brinkmann et al. (2023), repeated interactions with machines become increasingly important to maintain alignment Mikolov et al. (2018); Beuls and Van Eecke (2024) and referential grounding Kouwenhoven et al. (2022). In the case of humans, these repeated interactions cause languages to evolve such that they accommodate for specific abilities and preferences of minority individuals at the group level Josserand et al. (2024). Since the seemingly similar ways that languages adapt and optimise as a result of learning and use in Human-Human and LLM-LLM interactions, the question arises whether these processes can also be used to evolve a language that is optimised for humans and LLMs. In other words; can humans and LLMs collaboratively shape a language that is easy to learn for both, and allows for successful communication? If so, what do these languages look like?
@@
This work investigates these questions by extending earlier work by Kouwenhoven et al. (2025). Firstly, we provide experimental data of humans playing the same referential game that Kouwenhoven et al. (2025) used with LLMs, allowing comparisons between languages evolved through LLM-LLM interactions with those resulting from Human-Human interactions. Secondly, we experiment with human-LLM collaboration (Figure 1)111This study was approved by the ethics department of Leiden University (2024-03-11-R.E. de Kleijn-V1-5354). The code and data are available on OSF https://osf.io/ep6mw/. While it is unclear how human and LLM abilities exactly differ, this allows us to test whether an artificial holistic language can be optimised for the inductive biases of two different types of language learners. If shared vocabularies of signals and meanings emerge that allow for successful communication, one could argue that there has been some form of referential grounding, a prerequisite for successful communication Clark and Brennan (1991). Finally, Human-LLM collaboration allows investigating if and how the evolved language differs from languages that evolve within Human-Human and LLM-LLM interactions.
@@
Our results show that structured and referentially grounded languages can emerge when humans and LLMs interact repeatedly. The languages that emerge from these interactions tend to be more human-like than LLM-like, suggesting that the LLMs are flexible towards the strong human preferences that shape the languages. Finally, languages optimised for LLMs have less variation and are more degenerate than those optimised for humans.
@@ 2 Background
We discuss relevant research on language evolution, the role of inductive biases, and their relevance to LLM research.
@@ 2.1 Language evolution
Language allows us to communicate successfully because of the vocabulary we share, but also due to its open-ended nature, enabling the possibility to express novel meanings via compositional semantics. This defining feature of human language means that the meaning of any phrase is derived from the meanings of its individual components and the rules by which they are combined Hockett (1960). The evolution of compositionality has been investigated abundantly in the field of language evolution through human experiments (Kirby et al., 2008; Raviv et al., 2019, e.g.) and computational simulations (de Boer, 2006; Steels et al., 2012; Lazaridou and Baroni, 2020; Kouwenhoven et al., 2024, e.g.). These experiments typically involve learning artificial languages or playing a signalling game. Here, learning artificial languages imposes a constraint which is believed to lead to more structured languages Kirby et al. (2015). Communication in signalling games imposes a pressure for expressivity, requiring participants to develop a vocabulary of signal-meaning mappings that allows them to communicate about novel stimuli. In this case, some form of referential grounding must be established through the process of repeated interactions. Participants–human or machine–generally establish novel signal-meaning mappings quickly that enable successful communication.
@@ 2.2 Inductive biases
An important aspect in this paper is the notion of biases. Here, we do not so much focus on behavioural biases observed in humans (e.g., the confirmation bias), but are rather interested in implicit inductive biases that may result in biased language learning. This is relevant since seemingly arbitrary aspects of linguistic structure may actually result from general learning and processing biases deriving from the structure of thought processes, perceptuo-motor factors, cognitive limitations, and pragmatics Christiansen and Chater (2008). For humans, inductive biases like a preference for compressibility, simplicity, or efficiency Kirby et al. (2015); Tamariz and Kirby (2015); Gibson et al. (2019) naturally shape how languages evolve. Some even argue that humans’ cognitive limitations may be beneficial for language acquisition DeCaro et al. (2008); Poletiek et al. (2018). In the case of simulations of (reinforcement learning) agents, inductive biases typically do not match those present in humans. As such, they are often induced artificially by incorporating biases to guide learning dynamics as a means to recover human-like properties (Galke and Raviv, 2025, for a review see). In the case of LLMs, being fundamentally different from humans (although this is an ongoing debate Kozachkov et al. (2023)), we focus on increasingly apparent inductive biases of the transformer architecture Futrell and Mahowald (2025) that may influence how languages evolve in the context of our experiment.
@@
Although the underlying mechanisms of these biases differ between humans and machines, we find substantial overlap in terms of their behavioural effects. As such, we hypothesise that the aforementioned effects of continuous learning and use of language will also come into play when humans and machines collaborate and result in a language optimised for the preferences of both entities.
@@ 3 Methodology
This experiment revolves around the classical referential Lewis game as implemented in Kouwenhoven et al. (2025), who based their setup on previous work in emergent communication(Raviv et al., 2021; Kirby et al., 2015, e.g.). We extend the setup used by Kouwenhoven et al. (2025) to incorporate humans. In total, 45 participants participated in the experiment, 30 of whom formed 15 Human-Human pairs, and the remaining 15 interacted with an LLM in a Human-LLM setup. This allows us to directly compare languages adapted for human preferences to those adapted for the LLM-LLM simulations Kouwenhoven et al. (2025). But perhaps most interestingly, the Human-LLM condition provides an opportunity to investigate whether languages can be optimised for entities with different mechanisms and cognitive capacities (e.g., memory), and if so, we can see what these look like.
During the experiment, participants learn an artificial language and use it during communication. This language comprises a meaning space consisting of three attributes (shape, colour, and amount) that each can have three values, totalling to 27 unique stimuli. The corresponding labels are initialised following the design of Kirby et al. (2008), creating a holistic artificial language without structure (e.g., ”watopo”, ”sanasowi”, ”pikuku”) that contained a limited set of characters to prevent participants from writing English words. Participants first learn 15 random signal-meaning pairs individually through exposure, guessing, and labelling blocks. Hereafter, participants use the newly acquired language to communicate in a referential game. In this game, the speaker observes a target stimulus and labels it. Using this label as a signal, the listener is then tasked with identifying the correct target among three distractors. Cooperation is successful when the listeners’ guess matches the target. After the communication block, there is a testing block in which participants individually label 27 meanings, including 12 unseen meanings. The duration of the entire experiment is roughly 70 minutes. An experiment overview is provided in Figure 1.
@@ 3.1 LLMs as participants
Human participants learn the language by going through exposure and guessing blocks twice. They iteratively go over the 15 training stimuli and extract some apparent, but not present, patterns or consistencies. They are then tasked to label the stimuli, before moving on to the communication block in which they interact with another human or a LLM. In either case, they were told that they interacted with a human. The LLM agents, however, are not updated and receive instructions to learn the languages by virtue of in-context learning. Specifically, the stimuli are presented in a structured JSON-like format (Figure 2) that proved effective in Galke et al. (2024) and Kouwenhoven et al. (2025). As such, we assume that these signal-meaning mappings in the context of a prompt provide enough (distributional) information for a LLM to learn a mapping between the attributes of the stimuli and signal syllables (1). Although the prompt structure ‘invites’ the LLM to infer a signal from the stimulus attributes, we are agnostic about how exactly and what kind of mapping is deduced, but are interested in the resulting behaviours. In the experiments, we use the instruction-tuned variant of Llama 3 70B Llama Team (2024) with greedy sampling. We opt for an instruction-tuned model since this allows us to specify the need for communicative success. This potentially affects how the model’s inductive biases are expressed, but we leave this to future work.
@@
One of essentially two tasks is performed throughout the experiment: labelling or guessing. The labelling block and speaking in the communication block involve labelling, and the guessing block and discrimination during communication involve guessing. Generating signals is achieved through prompt completion. Guessing is done by prefilling the prompt with distractor stimuli or labels and selecting the item with the highest probability. This alleviates LLMs’ inconsistent behaviour in answering multiple-choice questions Khatun and Brown (2024), and follows recommendations from computational linguistics Hendrycks et al. (2021); Wang et al. (2024). During communication, we add a communicativeSuccess attribute set to if the previous interaction for this stimulus was successful and zero otherwise. This attribute functions as a memory between interactions and acts as a pressure for expressivity. In human language evolution, such pressure plays an important role since it prevents languages from becoming degenerate Kirby et al. (2015). Importantly, the agents observe the training vocabulary in their context with the current stimulus in the guessing and labelling block, rendering them as simple look-up tasks. We do, however, not include the current stimulus during communication and testing, requiring the agents to extract an appropriate mapping and generalise to new stimuli. Akin to common practice in older simulations (Steels et al., 2012, e.g.)), the agent vocabularies are updated when labels are generated after the labelling block and during the communication block. In particular, each time a label is generated for a stimulus, it replaces its current label value. Hence, the agent’s vocabularies of signal-meaning mappings evolve during the simulation. As such, prompts are slightly different after each interaction. Moreover, given the primacy and recency bias in LLMs Liu et al. (2024); Mina et al. (2025), we shuffle the vocabulary before creating prompts to account for unwanted ordering effects.
@@ 3.2 Metrics
Besides comparing the percentage of communicative success (PercCom), the primary goal of this work is to understand what a language looks like when optimised for different entities. Specifically, we investigate whether the languages display some degree of structure in the form of compositionality. In this experiment, this means that attribute values are denoted with label parts that are reused to describe other similar stimuli. Capturing this is not at all trivial, especially provided the freedom given to participants when they label stimuli. A common metric that gauges whether similar meanings map to similar signals is Topographic Similarity (Brighton and Kirby, 2006, TopSim). While providing a good indication of compositional language use, it does not take variability in language, such as word-order freedom, into account. It could therefore show an incomplete picture (i.e., a low TopSim) since languages can still be compositional despite having multiple word orders, the existence of synonyms, or homonyms Conklin and Smith (2023). Hence, we report multiple metrics in addition to TopSim that together indicate the degree of compositionality. Specifically, on synonymy (one-to-many mappings), homonymy (many-to-one), and word order freedom (Freedom), for which Conklin and Smith (2023) proposed entropy-based metrics. A language where each attribute value is encoded by a single character in a position has low entropy, and thus a low synonymy. Languages with a uniform distribution over all characters to refer to an attribute value have high synonymy. Homonymy is similar, it looks at how many attribute-values a character in a position can refer to, i.e., when characters can map to multiple attribute-values. Finally, we compute word-order freedom (freedom) to account for variability in the order by which labels are composed. It assesses whether each value of a specific attribute is encoded in a specific position of the label, i.e., there is little freedom, or whether attribute values can be encoded in any position of the label, i.e., displaying a high degree of word order freedom ().
@@
Systematic generalisation to novel stimuli is assessed through the generalisation score GenScore from Raviv et al. (2021). It gauges whether the labels produced for unknown (i.e., testing) stimuli are labelled in consistent ways with labels produced for similar known (i.e., training) stimuli. In addition to character-based metrics, we assess whether participants re-use parts of labels in different labels by computing the Ngram diversity Li et al. (2016) over all the produced labels in a block. Ngram diversity is the average ratio of unique vs. total Ngrams for in all labels. Low Ngram diversity implies that labels are composed of reused parts and high diversity means that labels do not share many Ngrams, thus are very different. The percentage of unique labels captures the degree of degeneracy (RatioUniLabels). Finally, we measure whether a pressure for communicative success, known to drive efficiency in human experiments Smith and Culbertson (2020), results in shorter labels using WordLength.